OpenStudy (anonymous):

Find the absolute extrema of the function y=-x^2+3x-5, on [-2,1]

OpenStudy (ghazi):

differentiate it twice, first

OpenStudy (anonymous):

find the vertex, check the second coordinate of the vertex and the endpoints

OpenStudy (anonymous):

don't differentiate at all

OpenStudy (anonymous):

why twice?

OpenStudy (ghazi):

@satellite73 it's easier

OpenStudy (anonymous):

well, my teacher wants us to do it with differentiating.

OpenStudy (anonymous):

vertex does not need calculus first coordinate of the vertex is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

fine, take the derivative, set it equal to zero you will still get \(-\frac{b}{2a}\)

OpenStudy (ghazi):

but calculus makes an oral calculation here , no, need to plot vertex...and see finally you used calculus

OpenStudy (anonymous):

second deriative tells you nothing, since it is simply \(2a\) a constant, only tells you it is a parabola that faces down, which you know by looking

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i used calc because it was requested. vertex is \((-\frac{b}{2a},f(-\frac{b}{2a}))\) calculus or no calculus

OpenStudy (anonymous):

in any case you have to check 2x+y=-11,3x-4y=11 \[f(-2),f(1), f(\frac{3}{2})\] largest is max, smallest is min

OpenStudy (anonymous):

y'=-2x+3

OpenStudy (ghazi):

okay then ...you can guide better, i could have explained what that constant -2 means but sir, you carry on :)

OpenStudy (anonymous):

i set it equal to 0 and got x=3/2

OpenStudy (anonymous):

yes of course as \(a=1,b=3,-\frac{b}{2a}=\frac{3}{2}\) so now you are left to check \[f(-2),f(1), f(\frac{3}{2})\]

OpenStudy (anonymous):

yeah, i got that. do i put it into y' or y?

OpenStudy (anonymous):

i.e. check the vertex and the endpoints of the interval to see which is largest and which is smallest

OpenStudy (anonymous):

so, y?

OpenStudy (anonymous):

you want the max and min of the function, put it in the function if you replace \(x\) by \(\frac{3}{12}\) i the derivative you will get zero, because that is how you found it to begin with

OpenStudy (anonymous):

yes use \(y\) not \(y'\) to evaluate

OpenStudy (anonymous):

so i got 3/2, -15, and -3. my answer sheet says -15 and -3 are the answer but i dont understand why.

OpenStudy (anonymous):

what is \(f(-2)\)?

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

oh wait, i forgot to put 3/2 into the eqwuation

OpenStudy (anonymous):

ok and \(f(1)=-3\) right?

OpenStudy (anonymous):

ignore the \(\frac{3}{2}\) it is not in your interval

OpenStudy (anonymous):

ohh! ok yeah i forgot about that. So how do I know which is the min and the max

OpenStudy (anonymous):

interval is \([-2,1]\) and your vertex is at \(x=1.5\) so forget it, function is going up on your interval

OpenStudy (anonymous):

not to be a wise guy, but the min is the min and the max is the max! you have \(-15\) at the left hand endpoint and \(-3\) at the right, evidently the min is \(-15\) and the max is \(-3\)

OpenStudy (anonymous):

that makes sense. thank you!

OpenStudy (anonymous):

don't forget minimum means minimum output, not input yw